Question on how to get this solution, see comments below and THX!

Effective cost of trade credit = [(1 + 0.0101)^(365 / 45)] - 1

                                                  = [(1.0101)^(365/45)] - 1

                                                 = [(1.0101)^8.11] - 1-------what is this symbol mean? multiple the two or how do I get 1.0849?

                                                 = 1.0849 -1

                                                 = 0.0849 or 8.49%
Answer #1

Explanation:

Effective cost of trade credit = [(1 + 0.0101) ^ (365 / 45)] - 1

                                                   = [(1.0101) ^ (365/45)] - 1

                                                   = [(1.0101) ^8.11] – 1

The symbol ^ means power to; for example 5^3 means 5 power to 3 or 5 multiplied three times in following manner

5^3 = 5 * 5 *5 = 125

Similarly [(1.0101) ^8.11] means 1.0101 power to 8.11

To calculate the value of [(1.0101) ^8.11] you can use calculator where you have to write 1.0101 then press symbol ^ and write 8.11, you will get 1.0849.

Or you can use excel in following manner

1.0101 8.11 1.0849

Therefore,

Effective cost of trade credit = 1.0849 -1

= 0.0849 (decimal form) to convert in in percentage for you have to multiply it with 100 therefore 0.0849 * 100 = 8.49%